OpenStudy (__meggaan):

For his chemistry experiment, Martin needs 3 liters of a 40% alcohol solution. The lab has two containers, one with 20% alcohol solution and the other with 50% alcohol solution. And I need to find a set of linear equations.

OpenStudy (dumbcow):

Let x be num of liters of 20% solution, y is num of liters of 50% solution -You need total of 3 liters of mixed 40% solution x + y = 3 -You can multiply the amount and the percentage to get alcohol content. - The final amount of alcohol must be 40% of 3 liters 0.2x + 0.5y = 0.4(3) Now you have a system of 2 equations which you can solve for x and y

OpenStudy (mhchen):

Use the equation M1V1 = M2V2